Take the following two exchange rates and compute the EUR/INR cross exchange rate. INR12.1225/USD and EUR.8145/USD....

  1. Take the following two exchange rates and compute the EUR/INR cross exchange rate. INR12.1225/USD and EUR.8145/USD.

  1. In question 4, if there is a direct cross exchange rate of EUR.066215/INR, is there a triangular arbitrage opportunity? If yes, start with $50,000 and indicate how much triangular arbitrage profit exists for 1 trip around the triangle.
Answer #1

HI,

Lets say we are having $50,000

with that you will but EUR based on exchange rate given.

1 USD = 0.8145 EUR

50000 USD = 50000*0.8145 = EUR 40,725

From that you will buy INR

0.066215 EUR =1 INR

1 EUR = 1/0.066215 INR

40725 EUR = 40725/0.066215 = INR 615,041.9089

From this INR again USD will be bought

12,1225 INR =1 USD

615,041.9089 INR = 615041.9089/12.1225 = $50,735.57

Hence arbitrage profit = 50735.57 - 50000 = $735.57

Thanks